Ahmer walks 200m east and 80m north 180m east find the displacement and distance?
1
Expert's answer
2020-03-05T09:35:45-0500

The distance


"d=200\:\rm m+80\:\rm m+180\:\rm m=460\:\rm m."

The displacement


"s=\sqrt{(200+180)^2+80^2}=388\:\rm m."